The perimeter of a rectangular field is 260 yards. If the length of the field is 78 yards, what is its width?
Answer by checkley77(12844)

2*78+2w=260
156+2w=260
2w=260-156
2w=104
w=104/2
w=52 yds. ans. for the width.
Proof:
2*78+2*52=260
156+2*52=260
156+104=260
260=260